# Modeling error correction with Lindblad dynamics and approximate channels

### v1

@@ IV.3 The Pauli approximation
The most precise way to perform the Pauli approximation is by approximating the full five-qubit quantum channel. But this will be less precise than applying the full channel, and will hardly, if at all, be advantageous in the ease of simulating. Here we choose a less exact but easier to compute approximation, which consists of approximating each individual channel separately. The channels $\mathcal{K}_{\text{1Q}}$ and $\mathcal{U}^{zz}_{\text{2Q}}$ are each approximated by 1Q and 2Q Pauli channels, respectively, which are then composed. A more detailed explanation on the different possible ways to perform this Pauli approximation is given in App. G. We note that the alternative method of approximating the 1Q and 2Q channels together into one 2Q Pauli channel is not a better choice due to the complication arising from the need to “split” the 1Q channels into the different 2Q channels in which each qubit is included, which reduces the accuracy of the approximation.
@@
It is important to note that the Pauli channels always commute with each other (App. E). Hence, for this approximation the order in which we apply the different channels does not matter.
@@
Finally, a natural question is whether a 1Q approximation can be devised that is reliable for capturing the 2Q noise, as depicted in Figure 1(c). Such an approximation – and in particular if it could be reduced to a 1Q Pauli noise channel – could be very useful for error correction threshold estimations. This would require approximating $\mathcal{U}^{zz}_{\text{2Q}}$ by a single qubit channel, since the other channels discussed above are already 1Q channels. When a two-qubit unitary acts on a separable state $|\psi_{1}\psi_{2}\rangle$, the effect of this on one of the qubits – disregarding the second – is described by a channel with the Kraus operators $E_{k}=\left\langle e_{k}|U|\psi_{2}\right\rangle$ where $\left\{|e_{k}\rangle\right\}$ is a basis for the first qubit. However, this approach isn’t applicable in our case due to the highly entangled nature of the logical states. We tried to create a Kraus channel when assuming the qubit is not-entangled to its environment, or by trying to capture the basic actions of the channel on each qubit, but did not manage to create a sufficiently adequate 1Q approximation. As presented in more detail in App. J, we could not find such an approximation that is controlled and captures the 2Q noise dynamics in a satisfactory way. This implies that whenever the 2Q noise magnitude is at least comparable to the other noise terms, it has to be properly accounted for and a 1Q approximation cannot be used.
@@ V Results
In this section we present our main results and highlight the cases where certain approximations hold and which noise terms cause them to break down. We first note that with 1Q noise only, i.e., for $\zeta=0$, all noise channels commute and so the composite-channel approximation coincides with the dynamical simulation. Examining the Pauli approximation for different timescales and different initial states, we found it to be very accurate also, and a few examples are presented in App. J. We therefore focus for the rest of this section on the regime of nonnegligible 2Q noise.
For completeness, we note that the logical code states of the five-qubit code that we investigate are defined as the $+1$ eigenstates of all four stabilizers
$S_{1}$ $=X_{0}Z_{1}Z_{2}X_{3}I_{4},$ (38)
$S_{2}$ $=I_{0}X_{1}Z_{2}Z_{3}X_{4},$ (39)
$S_{3}$ $=X_{0}I_{1}X_{2}Z_{3}Z_{4},$ (40)
$S_{4}$ $=Z_{0}X_{1}I_{2}X_{3}Z_{4},$ (41)
where $\{X,Y,Z\}$ denote the three Pauli matrices, and the indices indicate the qubit. The logical Pauli operators are
$Z_{L}$ $=Z_{0}Z_{1}Z_{2}Z_{3}Z_{4},$ (42)
$X_{L}$ $=X_{0}X_{1}X_{2}X_{3}X_{4}.$ (43)
Dynamical timescales and parameter regimes
Before presenting our results in the following subsections, we discuss the relevant parameter regimes that this section focuses on. As mentioned in Sec. II, we consider parameters that are identical for all qubits. The values of the noise parameters are taken to be representative of realistic parameters of superconducting-qubit devices, as presented in App. A. We first note that the 2Q noise term cannot be neglected with existing superconducting qubits, as can be seen by examining the relevant nondimensional ratios $|\zeta/h|$ and $|\zeta|T_{1}$. Substituting $h=2\pi\times 5\,$kHz, $\zeta=2\pi\times 3\,$kHz and $T_{1}=150\,\mu$s as typical values obtained for a device of Heron type, we find that the qubits of this device are typically in the regimes of $|\zeta/h|\sim 1$ and $|\zeta|T_{1}\gtrsim 1$, where the coherent 1Q and 2Q terms are comparable and larger than the incoherent noise. This obviously holds also in typical Eagle devices where $\zeta$ is larger by an order of magnitude. Qubit coherence times are constantly being improved, and with $T_{1}=1$ms, the crosstalk of idle qubits would have to be significantly reduced down to $\zeta=2\pi\times 0.1\,$kHz in order to change into the regime of $|\zeta|T_{1}\lesssim 1$.
Based on this observation, we arrange our results according to three dynamical timescales. We denote these as the short, intermediate, and long times. The long time limit is defined to be that in which $\rho(t)$ approaches a steady state, for $t\gg T_{1}$. We consider the timescale $t_{\rm L}\sim T_{1}$ as the border of the long time limit, which for concreteness can be taken to be at $100\mu$s. This limit is explored briefly in subsection V.1, and is obviously outside of the regime of useful error correction.
The intermediate timescale is set by the largest error term that we consider, and can be defined by as the time $t_{\rm I}$ obeying $|\zeta|t_{\rm I}\sim 1$. For concreteness, that would be $5\,\mu{\rm s}$ for the typical Eagle value $\zeta=-2\pi\times 30\,$kHz, and it would be $50\,\mu{\rm s}$ with the Heron value of $\zeta=2\pi\times 3\,$kHz.
@@
The timescales described above and their definitions are summarized in Table 1. In the following subsections we arrange the results according to the short and intermediate timescales, and analyze the times at which the various approximations break down quantitatively as well as qualitatively.
@@
V.1 The long time limit
The long-time limit is not relevant for error correction, which is applicable when the error rate is low. Nevertheless, it is informative since it manifests a fundamental difference between the Pauli approximation to the composite-channel and Lindblad dynamics.
V.2 Dominant ZZ crosstalk noise
Figure 2: The failure probability after error recovery – $\eta$ of Eq. (10) – vs. $t$, for $t$ the time of the noise channel, and the noise parameters in Eqs. (44)-(45), with dominant crosstalk noise. (a) An average over the initial states in Eq. (11), in the short-time regime, showing that all approximations are identical. (b) The initial state is $|0\rangle_{L}$, shown up to the intermediate timescale, when the Pauli approximation breaks, and also the composite-channel approximation loses accuracy (shown for two possible orderings, see the text for details).
@@ V.2 Dominant ZZ crosstalk noise
Table 2: Table of values of $\eta$ for different approximations and different noise durations. The noise parameters are given in Eqs. (44)-(45). The values presented in this table are $\left\langle\eta\right\rangle\left(\text{std}\left(\eta\right)\right),\sqrt{\left\langle\left(\eta-\eta_{\rm{D}}\right)^{2}\right\rangle}$ where $\eta_{\rm{D}}$ is $\eta$ under the dynamical simulation of the noise, and the average and standard deviation are taken over the six Pauli eigenstates as initial states [Eq. (11)]. In this case the composite-channel approximation does very well, while the Pauli approximation is good for short time scales, but differs significantly for longer time-scales.
@@
In Fig. 2(a) we show the average of $\eta$ over the initial states in Eq. (11), in the short-time regime. Since the separate channels in the composite approximation do not commute, the order in which they are applied matters. In Fig. 2 composite 1 refers to applying the 1Q channels before the 2Q channels, and composite 2 is the reverse order. The results are very similar for every initial state separately. It can be seen that all approximations do very well. In Fig. 2(b), we examine times up to the intermediate timescale, focusing on the initial state $|0\rangle_{L}$, for which the Pauli approximation differs the most. Here we can see that the Pauli approximation differs significantly and qualitatively, but for long-enough times, also the composite-channel approximation deviates.
Table 2 compares the accuracy of the different approximations when averaged over the initial states in Eq. (11). In the table we present the expectation value and standard deviation of $\eta$. In order to express the variance of the success rate for the initial state, we also present the expectation, over initial logical states, of the difference between $\eta$ of the dynamical simulation and that of the approximation.
@@
Regarding the failure of the Pauli approximation, there is a more complex mechanism at play here. In order to gain more insight into it, we consider in the next subsection the question of a decoder tailored to the noise.
V.3 Modified decoders
The five-qubit code with the standard decoder is a distance-three code, i.e., it can successfully correct any single 1Q error. However, since the noise model includes also 2Q ZZ terms, when the probability of those 2Q errors is larger than that of (at least some of) the 1Q errors, it could be beneficial to sacrifice the correction of 1Q errors in order to gain the ability to correct the more prevalent 2Q errors. In this case the decoder can be best constructed to account also for qubit connectivity. In App. K we present in detail how the standard five-qubit code decoder can be modified in order to correct one $Z_{i}Z_{j}$ error on any pairs of qubits. This is achieved without changing the code states or syndromes but rather reinterpreting them, by sacrificing the ability to correct single qubit $X$ or $Y$ errors. We denote this modified decoder as the full-connectivity (FC) ZZ decoder.
In Fig. 3 we present the difference between $\eta$ with the FC ZZ decoder and with the standard decoder, showing that the FC ZZ decoder is beneficial in a wide range of experimentally relevant parameters. Specifically, the Eagle device parameters [Eqs. (44)-(45)] fall in the regime where the modified decoder is advantageous. Therefore, applying the FC ZZ decoder to the dynamics with the same parameters as in Fig. 2(a), the FC ZZ decoder performs much better than the regular decoder, obtaining much lower values of $\eta$, as can be seen in Fig. 4. We also notice that the Pauli approximation begins to diverge from the dynamical simulation as early as $t\approx 0.7\mu$s.
In App. K we present in more detail results for the case of qubits in a ring topology, using an adequate choice of a slightly different decoder. We also show that with a decoder that does not match the qubits’ connectivity, the failure rate is increased.
@@ V.3 Modified decoders
V.4 Purely 2Q ZZ crosstalk
Figure 5: The failure probability $\eta$ vs. $|\zeta|t$, for $t$ the time of the noise channel, with only ZZ crosstalk noise, of magnitude $|\zeta|$ [Eq. (46)]. The composite-channel approximation is exact and identical to the full dynamical solution in this case. The curves are an average over the six logical Pauli eigenstates [Eq. (11)] in the short-time limit, with (a) the regular decoder, and (b) the FC ZZ decoder. We can see that the latter decoder outperforms the former (by about an order of magnitude). The Pauli approximation holds until $|\zeta|t\approx 0.2$ for the regular decoder, while it doesn’t hold for any value of $|\zeta|t$ for the FC ZZ decoder.
Noticing that the 2Q noise term is the most dominant in the Eagle device we are interested in analyzing the case where the sole source of noise is the 2Q term $V^{zz}$. And so we consider the case with parameters
@@ V.4 Purely 2Q ZZ crosstalk
In this case, since the noise terms on all pairs of qubits commute, the composite-channel approximation is equivalent to the dynamical simulation, and our focus is on the validity of the Pauli approximation. The errors only depend on the combination $\zeta t$, and in the current case with $h=0$, we observe the dynamics are independent of the sign of $\zeta$.
For the regular decoder, we see in Fig. 5(a) that the Pauli approximation does very well up to $|\zeta|t\approx 0.2$ on the average of the initial states. We note that there is a vast difference between initial states, with the approximation holding for much larger values for some initial states. The breaking points of the approximation are given in terms of the variable $|\zeta|t$, implying that the time $t$ of the approximation breaking point would be very different for different values of $\zeta$. For example, the value $|\zeta|t=0.2$ is obtained at $t\approx 1\,\mu$s with $\zeta=-2\pi\times 30\,$kHz, while for $\zeta=2\pi\times 3\,$kHz it is reached at $t\approx 10\,\mu$s.
The FC ZZ decoder, presented in Fig. 5(b) outperforms the regular decoder for small values of $|\zeta|t$, as expected. It is interesting to note that for this decoder, the Pauli approximation does not hold for any value of $|\zeta|t$, no matter how small.
V.5 Lower crosstalk noise
@@
and for the dissipative parameter we keep as before,
@@
V.6 Nonsymmetric reference frame
Figure 7: The failure probability $\eta$ vs. $t$, where the noise parameters are $\tilde{h}=2\pi\times 5$ kHz, $\zeta=2\pi\times 3$ kHz, $T_{1}=150$ $\mu$s and $T_{2}=100$ $\mu$s, in the short time scale. The initial states are, (a) $|0\rangle_{L}$, (b) $|+\rangle_{L}$. Performing the Pauli approximation via adjusting the 1Q term is more accurate in some cases, but not all.

### v2

@@ IV.3 The Pauli approximation
The most precise way to perform the Pauli approximation is by approximating the whole-system quantum channel. However, this requires constructing the channel which may be numerically as challenging as simulating its action on particular state (or even more). Here we choose a less exact but easier to compute approximation, which consists of approximating each individual channel separately. The channels $\mathcal{K}_{\text{1Q}}$ and $\mathcal{U}^{zz}_{\text{2Q}}$ are each approximated by 1Q and 2Q Pauli channels, respectively, which are then composed. A more detailed explanation on the different possible ways to perform this Pauli approximation is given in App. G. We note that the alternative method of approximating the 1Q and 2Q channels together into one 2Q Pauli channel is not a better choice due to the complication arising from the need to “split” the 1Q channels into the different 2Q channels in which each qubit is included, which reduces the accuracy of the approximation.
@@
Since the Pauli channels always commute with each other (App. E), for this approximation the order in which we apply the different channels does not matter.
@@
Finally, a natural question is whether a 1Q approximation can be devised that is reliable for capturing the 2Q noise, as depicted in Figure 1(c). Such an approximation – and in particular if it could be reduced to a 1Q Pauli noise channel – could be very useful for error correction threshold estimations. This would require approximating $\mathcal{U}^{zz}_{\text{2Q}}$ by a single-qubit channel, since the other channels discussed above are already 1Q channels. When a two-qubit unitary acts on a separable state $|\psi_{1}\psi_{2}\rangle$, the effect of this on one of the qubits – disregarding the second – is described by a channel with the Kraus operators $E_{k}=\left\langle e_{k}|U|\psi_{2}\right\rangle$ where $\left\{|e_{k}\rangle\right\}$ is a basis for the first qubit. However, this approach isn’t applicable in our case due to the highly entangled nature of the logical states.
We tried to create an approximated 1Q noise channel by approximating each qubit to be separable from its neighbors and applying $E_{k}=\left\langle e_{k}|U|\psi_{2}\right\rangle$, or by identifying the basic actions of the 2Q noise on each qubit, and building a 1Q channel with similar properties. The 1Q channels we constructed differed qualitatively from the code behavior under 2Q noise at all timescales, and weren’t sufficiently adequate even for very short times. The details of the 1Q approximations we constructed, and their inferior performance are presented in App. J. This implies that whenever the 2Q noise magnitude is at least comparable to the other noise terms, it has to be properly accounted for and a 1Q approximation cannot be used.
@@ V Results
In this section we present our main results and highlight the cases where certain approximations hold and what are the mechanisms of their breaking.
V.1 Dynamical timescales and parameter regimes
We first note that with 1Q noise only, i.e., for $\zeta=0$, all noise channels commute and so the composite-channel approximation coincides with the dynamical simulation. Examining the Pauli approximation for different timescales and different initial states, we found it to be very accurate also, and a few examples are presented in App. J. As mentioned in Sec. II, we consider parameters that are identical for all qubits, and taken to be representative of realistic parameters of superconducting-qubit devices, as presented in App. A. Indeed, the 2Q noise term of idle qubits cannot be neglected with existing superconducting qubits, as can be seen by examining the relevant nondimensional ratios $|\zeta/h|$ and $|\zeta|T_{1}$. Substituting $h=2\pi\times 5\,$kHz, $\zeta=2\pi\times 3\,$kHz and $T_{1}=150\,\mu$s as typical values obtained for a device of Heron type, we find that the qubits of this device are typically in the regimes of $|\zeta/h|\sim 1$ and $|\zeta|T_{1}\gtrsim 1$, where the coherent 1Q and 2Q terms are comparable and larger than the incoherent noise. This obviously holds also in typical Eagle devices where $\zeta$ is larger by an order of magnitude. We therefore focus for the rest of this section on the regime of nonnegligible 2Q noise.
Based on these observations, we arrange our results according to three dynamical timescales. We denote these as the short, intermediate, and long times. The long time limit is defined to be that in which $\rho(t)$ approaches a steady state, for $t\gg T_{1}$. We consider the timescale $t_{\rm L}\sim T_{1}$ as the border of the long time limit, which for concreteness can be taken to be at $100\mu$s. This limit is explored briefly in subsection V.2, and is obviously outside of the regime of useful error correction.
The intermediate timescale is set by the largest error term that we consider, and can be defined as the time $t_{\rm I}$ obeying $|\zeta|t_{\rm I}\sim 1$. For concreteness, that would be $5\,\mu{\rm s}$ for the typical Eagle value $\zeta=-2\pi\times 30\,$kHz, and it would be $50\,\mu{\rm s}$ with the Heron value of $\zeta=2\pi\times 3\,$kHz.
@@ V.1 Dynamical timescales and parameter regimes
The timescales described above and their definitions are summarized in Table 1. In the following subsections we analyze the times at which the various approximations break down quantitatively as well as qualitatively.
@@
V.2 The long time limit
The long-time limit of a noise channel and a single syndrome cycle is not relevant for error correction. Nevertheless, it is informative since it manifests a fundamental difference between the Pauli approximation to the composite-channel and Lindblad dynamics.
@@ V.2 The long time limit
Table 2: Table of values of $\eta$ for different approximations and different noise durations. The noise parameters are given in Eqs. (40)-(41). The values presented in this table are $\left\langle\eta\right\rangle\left(\text{std}\left(\eta\right)\right),\sqrt{\left\langle\left(\eta-\eta_{\rm{D}}\right)^{2}\right\rangle}$ where $\eta_{\rm{D}}$ is $\eta$ under the dynamical simulation of the noise, and the average and standard deviation are taken over the six Pauli eigenstates as initial states [Eq. (9)]. In this case the composite-channel approximation does very well, while the Pauli approximation is good for short time scales, but differs significantly for longer time-scales, for which, in this example, it overestimates the failure rate.
V.3 Dominant ZZ crosstalk noise
@@
In Fig. 2(a) we show the average of $\eta$ over the initial states in Eq. (9), in the short-time regime. Since the separate channels in the composite approximation do not commute, the order in which they are applied matters. In Fig. 2, the label “composite 1” refers to applying the 1Q channels before the 2Q channels, and “composite 2” implies the reverse order. The results are very similar for every initial state separately. It can be seen that all approximations do very well. In Fig. 2(b), we examine times up to the intermediate timescale, focusing on the initial state $|0\rangle_{L}$, for which the Pauli approximation differs significantly and qualitatively, but for long-enough times, also the composite-channel approximation deviates.
Figure 2: The failure probability after error recovery – $\eta$ of Eq. (11) – vs. $t$, the duration of the noise channel, and the noise parameters in Eqs. (40)-(41), with dominant crosstalk noise. (a) An average over the initial states in Eq. (9), in the short-time regime, showing that all approximations are identical. (b) The initial state is $|0\rangle_{L}$, shown up to the intermediate timescale, when the Pauli approximation breaks, and also the composite-channel approximation loses accuracy (shown for two possible orderings, see the text for details).
Table 2 compares the accuracy of the different approximations, by showing the mean and standard deviation when averaged over the initial states in Eq. (9). In order to express the dependence of the success rate on the initial state, we also present the standard deviation over initial logical states, of the difference between $\eta$ of the dynamical simulation and that of the approximation. We note that the pseudo-threshold in this case is very low (App. C) since the noise is dominantly 2Q noise and when being compared with a physical (single) qubit, the single qubit does not suffer from this noise.
@@
In the current example, the crosstalk noise is the largest term, and is one that cannot be corrected with the employed standard decoder of the 5Q code. The Pauli approximation overestimates the failure rate when leaving the short time regime, plausibly since it does not describe the coherent dynamics accurately enough. In order to gain more insight into it, we consider in the next subsection the question of a decoder tailored to the noise.
V.4 Modified decoders
The five-qubit code with the standard decoder is a distance-three code, i.e., it can successfully correct any single 1Q error. However, since the noise model includes also 2Q ZZ terms, when the probability of those 2Q errors is larger than that of (at least some of) the 1Q errors, it could be beneficial to sacrifice the correction of 1Q errors in order to gain the ability to correct the more prevalent 2Q errors. In this case the decoder can be best constructed to account also for qubit connectivity. In App. L we present in detail how the standard five-qubit code decoder can be modified in order to correct one $Z_{i}Z_{j}$ error on any pairs of qubits. This is achieved without changing the code states or syndromes but rather reinterpreting them, by sacrificing the ability to correct single qubit $X$ or $Y$ errors (while keeping the correction of all $Z$ errors untouched). We denote this modified decoder as the full-connectivity (FC) ZZ decoder.
In Fig. 3 we present the difference between $\eta$ with the FC ZZ decoder and with the standard decoder, showing that the FC ZZ decoder is beneficial in a wide range of experimentally relevant parameters. Specifically, the parameter values of Eqs. (40)-(41) fall in the regime where the modified decoder is advantageous. Therefore, applying the FC ZZ decoder to the dynamics with the same parameters as in Fig. 2(a), the FC ZZ decoder performs much better than the regular decoder, obtaining much lower values of $\eta$, as can be seen in Fig. 4. We also notice that the Pauli approximation begins to diverge from the dynamical simulation as early as $t\approx 0.7\mu$s, and now it underestimates the failure rate. In other words, even at lower failure rates and when the code is adapted to the dominant noise terms, the Pauli approximation could lose its accuracy already in the short time regime, and not provide even an upper bound on the logical error.
In App. L we present in more detail results for the case of qubits in a ring topology, using an adequate choice of a slightly different decoder. We also show that with a decoder that does not match the qubits’ connectivity, the failure rate is increased.
@@ V.4 Modified decoders
V.5 Purely ZZ crosstalk
Figure 5: The failure probability $\eta$ vs. $|\zeta|t$, for $t$ the time of the noise channel, with only ZZ crosstalk noise, of magnitude $|\zeta|$ [Eq. (42)]. The composite-channel approximation is exact and identical to the full dynamical solution in this case. The curves are an average over the six logical Pauli eigenstates [Eq. (9)] in the short-time limit, with (a) the regular decoder, and (b) the FC ZZ decoder (the inset showing the same data using log-log axes). We can see that the latter decoder outperforms the former (by about an order of magnitude), as expected. The Pauli approximation holds until $|\zeta|t\approx 0.2$ for the regular decoder, while it doesn’t hold for any value of $|\zeta|t$ for the FC ZZ decoder. See the text for a detailed discussion.
In order to gain a deeper understanding of the Pauli approximation failure we turn to analyze the case where the sole source of noise is the 2Q term $V^{zz}$, i.e., we set
@@ V.5 Purely ZZ crosstalk
In this case, since the noise terms on all pairs of qubits commute, the composite-channel approximation is equivalent to the dynamical simulation, and, as stated above, our focus is on the validity of the Pauli approximation. The errors only depend on the combination $\zeta t$, and in the current case with $h=0$, we observe that the dynamics are independent of the sign of $\zeta$.
For the regular decoder, we see in Fig. 5(a) that the Pauli approximation does very well up to $|\zeta|t\approx 0.2$ on the average of the initial states. For concreteness, the value $|\zeta|t=0.2$ is obtained at $t\approx 1\,\mu$s with $\zeta=-2\pi\times 30\,$kHz, while for $\zeta=2\pi\times 3\,$kHz it is reached at $t\approx 10\,\mu$s. We note that there are large differences between initial states, with the approximation holding for much larger values for some initial states.
The FC ZZ decoder, presented in Fig. 5(b), outperforms the regular decoder for small values of $|\zeta|t$, as expected and is evident by the failure rates smaller by an order of magnitude. However, it can be seen that for this decoder, the Pauli approximation does not hold for any value of $|\zeta|t$. An expansion analysis in fact shows that in this case, the Pauli approximation constructed from the 2Q noise parameters (and corresponding generators) underestimates the error by a factor of 3 (up to perturbative corrections) for some states and any noise amplitude $\zeta$. This factor is derived in App. K where it is seen to stem from unitary evolution terms in second order that are neglected in the construction of the Pauli approximation. As discussed in App. K and also further in Sec. V.8, this scaling factor of the crosstalk terms in the Pauli channel in the general case depends on the details of other noise terms (and the particular state, and decoder), and therefore needs to be derived numerically by comparison of the failure rate to the dynamics.
V.6 Lower crosstalk noise
@@
and for the dissipative parameters we keep as before,
@@
V.7 Nonsymmetric reference frame
Figure 7: The failure probability $\eta$ vs. $t$, where the noise parameters are $\tilde{h}=2\pi\times 5$ kHz, $\zeta=2\pi\times 3$ kHz, $T_{1}=150$ $\mu$s and $T_{2}=100$ $\mu$s, in the short time scale. The initial states are: (a) $|+\rangle_{L}$, (b) $|0\rangle_{L}$. Performing the Pauli approximation via adjusting the 1Q term could be more accurate in some cases but is not decisively so, and it might both overestimate and underestimate the failure rate, already in the short time regime.